QUASICONFORMALITY AND GEOMETRICAL FINITENESS IN

CARNOT–CARATHÉODORY AND NEGATIVELY CURVED SPACES

BORIS APANASOV

Abstract. The paper sketches a recent progress and formulates several open problems in studying equivariant quasiconformal and quasisymmetric homeomorphisms in negatively curved spaces as well as geometry and topology of noncompact geometrically finite negatively curved manifolds and their boundaries at infinity having Carnot–Carathéodory structures. Especially, the most interesting are complex hyperbolic manifolds with Cauchy–Riemannian structure at infinity, which occupy a distinguished niche and whose properties make them surprisingly different from real hyperbolic ones.